What is the angle to the third-order dark fringe (m = 3) for a single slit of width 37×10−5m if the wavelength of light is 582 nm?
Answer #1

For dark fringe , d Sin(theta) = m(lemda)

=) Sin(theta) = 3×582 × 10^-9 / 37×10^-5 = 47.189×10^-4

=) Theta = Sin^-1(0.0047189) = 0.27 degree

Hence angle theta = 0.27 °